OpenStudy (anonymous):

simplify: sin(sec^-1 x+1/3) I have the answer, just not sure how to get there.

OpenStudy (xapproachesinfinity):

you just asked a similar question earlier didn't you? it the same process

OpenStudy (anonymous):

what is secant though

OpenStudy (xapproachesinfinity):

first is that \(\large \rm sin(sec^{-1}(x)+1/3)\)

OpenStudy (xapproachesinfinity):

secx is just 1/cosx

OpenStudy (xapproachesinfinity):

is what i wrote correct?

OpenStudy (anonymous):

yes the answer is square root x^2 + 2x -8 all over x+1 but where did the polynomial come from

OpenStudy (xapproachesinfinity):

I'm asking did you write your problem like this \(\large \rm sin(sec^{-1}(x+1/3))\) or \(\large \rm sin(sec^{-1}(x)+1/3)\) i'm less concerned about the answer they give you make sure you wrote the right thing.Also, make sure you looking for the right question

OpenStudy (anonymous):

I meant it like the first way sorry about that

OpenStudy (xapproachesinfinity):

okay first set \(\large \rm sec^{-1}(x+1/3)=\theta\) so \(\large \rm sin\theta=x+1/3\)

OpenStudy (xapproachesinfinity):

let's draw a picture that goes with this information

OpenStudy (xapproachesinfinity):

im sorry should be sec(theta)=x+1/3 not sin

OpenStudy (xapproachesinfinity):

 we need to find the missing side

OpenStudy (xapproachesinfinity):

ops sorry, forgot what i did! the first one you said was the right expression i was working on the second one lol

OpenStudy (xapproachesinfinity):

let's rebuild! \(\large \rm sec^{-1}(x)=\theta ~~\Longrightarrow sec\theta=x\)

OpenStudy (xapproachesinfinity):

I will need to ask you to recheck your problem again! and rewrite the exact expression alright! use the drawing tool